### v1

# ADS-Edit: A Multimodal Knowledge Editing Dataset for Autonomous Driving Systems

Abstract
Recent advancements in Large Multimodal Models (LMMs) have shown promise in Autonomous Driving Systems (ADS). However, their direct application to ADS is hindered by challenges such as misunderstanding of traffic knowledge, complex road conditions, and diverse states of vehicle. To address these challenges, we propose the use of Knowledge Editing, which enables targeted modifications to a model’s behavior without the need for full retraining. Meanwhile, we introduce ADS-Edit, a multimodal knowledge editing dataset specifically designed for ADS, which includes various real-world scenarios, multiple data types, and comprehensive evaluation metrics. We conduct comprehensive experiments and derive several interesting conclusions. We hope that our work will contribute to the further advancement of knowledge editing applications in the field of autonomous driving111Code and data are available in https://github.com/zjunlp/EasyEdit..
ADS-Edit: A Multimodal Knowledge Editing Dataset for
Autonomous Driving Systems
Chenxi Wang♡11footnotemark: 1, Jizhan Fang♡11footnotemark: 1, Xiang Chen♠††thanks: Equal Contribution., Bozhong Tian♡, Ziwen Xu♡, Huajun Chen♡, Ningyu Zhang♡††thanks: Corresponding author. ♡Zhejiang University ♠Nanjing University of Aeronautics and Astronautics {sunnywcx,zhangningyu}@zju.edu.cn
1 Introduction
The recent Large Multimodal Models (LMMs) (Wang et al., 2024a; Li et al., 2024a; Wu et al., 2024b; Liu et al., 2024c; Team, 2024, 2025) have significantly enhanced capabilities in video understanding and multimodal reasoning. As a key foundation, LMMs have also found initial applications in Autonomous Driving Systems (ADS) (Xing et al., 2024; Sima et al., 2024; Huang et al., 2024b; Yao et al., 2024). However, the direct application of LMMs in ADS yields suboptimal results.
As shown in Figure 1, the reasons for failure can be attributed to the following factors: 1) Traffic knowledge misunderstanding: General models misunderstand traffic knowledge leading to suboptimal performance in tasks. (2) Complex and varied road condition: Real-world driving scenarios are highly variable, with training datasets often failing to cover edge cases. (3) Diversity of vehicle motion states: Current LMMs struggle to predict unknown and highly dynamic vehicle motion states. These challenges require the model to have the capability of updating knowledge in real-time and continuously.
To address these challenges, we propose the use of Knowledge Editing, which enables targeted modifications to the model’s behavior without the computational burden of full retraining. Unlike traditional fine-tuning, which risks catastrophic forgetting and demands extensive resources, Knowledge Editing facilitate rapid knowledge updates by selectively altering parameters associated with specific factual or contextual knowledge (Yao et al., 2023; Gupta et al., 2024; Zhang et al., 2024d; Fang et al., 2024a; Youssef et al., 2025; Jiang et al., 2025; Liu et al., 2024a; Yao et al., 2025). Although Knowledge Editing has been preliminarily explored in the area of LMMs, existing efforts have primarily focused on editing multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024) and multimodal factual knowledge (Zhang et al., 2024a; Huang et al., 2024a). To address this limitation, we pioneer the editing of multimodal domain-specific knowledge in the field of autonomous driving.
Following the principle of leveraging knowledge editing to address the challenges of autonomous driving, we design the ADS-Edit benchmark. This benchmark encompasses three real-world scenarios: perception, understanding, and decision making. Additionally, it incorporates three types of data: video, multi-view images, and single image. Furthermore, we establish comprehensive evaluation metrics for knowledge editing in autonomous driving scenarios.
We evaluate four commonly used Knowledge Editing baselines under both single editing and lifelong editing scenarios, such as Prompt, AdaLora (Zhang et al., 2023), GRACE (Hartvigsen et al., 2023), and WISE (Wang et al., 2024b). Thirdly, through our analysis, we obtained a series of interesting findings, including the universality of knowledge editing methods in updating knowledge across various scenarios, their ability to balance editing effectiveness and processing speed, particularly in the context of video data, and the remaining limitations in locality. In general, we conclude our contributions as:
We are the first to attempt Knowledge Editing on multimodal domain knowledge data, specifically the ADS and effectively addresses the current challenges faced by LMMs when directly applied to ADS.
We unveil ADS-Edit, a Knowledge Editing dataset specifically designed for the ADS. The dataset includes three common types of visual data and encompasses data that evaluates various model capabilities.
We test four Knowledge Editing baselines under both single editing and lifelong editing settings and analyze several interesting results.
2 Background
LMMs have been applied in autonomous driving, typically after being trained on carefully curated driving datasets. However, when faced with unfamiliar driving scenarios, such as new traffic regulations or predicting driver behavior during traffic congestion, the reliability of LMMs’ decision making can significantly degrade. Furthermore, LMMs struggle to maintain consistent performance when encountering sudden changes in road conditions, such as shifts in weather or unexpected traffic accidents. Finally, for continuously collected driving data, the challenge of updating knowledge in a timely and effective manner remains unresolved for LMMs. Overall, there is a critical need for a framework that enables LMMs to rapidly and sequentially update knowledges when applied to autonomous driving.
We propose leveraging knowledge editing techniques to address challenges in autonomous driving scenarios, defined as follows: Let be a LMM and be an edited model. Given the user’s inputs (Includes text and autonomous driving’s multimodal inputs , such as image or videos) and the editing target , the edited model is expected to modify LMM’s outputs within the editing scope to match the editing target, while preserving the original model’s output for the inputs outside the editing scope. The specific definition is as follows:
where and means the in-scope of the editing inputs.
3 Bechmark Construction
3.1 Desgin Principle
To construct a comprehensive benchmark, we propose a tri-axis design principle. This principle organizes evaluation requirements into scenario types: perception, understanding, and decision making, which progressively assess LMMs’ capabilities from basic visual recognition to complex behavioral reasoning. Concurrently, we categorize input modalities into data types: video, multi-view images, and single image. Multiple metrics are designed to evaluate knowledge editing methods, such as reliability, generality and locality.
Scenario Type. Perception scenario evaluates LMMs’ basic visual perception capabilities, such as obstacle detection and vehicle recognition. Understanding scenario requires the model to comprehend domain-specific knowledge of autonomous driving, such as traffic rules, beyond basic perception capabilities. Decision Making scenario presents a greater challenge, as it requires the model to integrate perception and understanding capabilities to make informed decisions about future driving behaviors. The statistics of scenario types is shown in Figure 2.
Data Type.
Video requires the model to possess the ability to evaluate temporal changes in images. Specifically, the model must be capable of understanding the three dimensions of image width, height, and time. Multi-view images are designed to assess the model’s capability when provided with sensor images from multiple viewpoints on the vehicle. Single image primarily tests the model’s fundamental perception abilities, such as object recognition and spatial relationship understanding. We will organize these diverse data types to construct a comprehensive benchmark. The statistics of different types of data is shown in Table 1.
| Video | Multi-view | Single | All | |
| Train | 1,926 | 960 | 1,093 | 3,979 |
| Test | 481 | 239 | 358 | 1,078 |
Metrics.
Reliability is to evaluate the success of behavioral modification in the target driving scenarios. Generality evaluates the generalization scores when facing similar autonomous driving scenarios. Locality measures whether methods alter unrelated knowledge after updating with domain-specific knowledge of autonomous driving.
3.2 Data Collection
Inspired by knowledge editing applications in unimodal settings, the constructed ADS-Edit dataset is composed of visual question answering data. The construction process of the ADS dataset is illustrated in Figure 3.
3.2.1 Reliability Data Construction
We select three well-known datasets of the autonomous driving system: LingoQA (Marcu et al., 2024), DriveLM dataset (Sima et al., 2024), and CODA-LM dataset (Li et al., 2024d) as raw data. Their autonomous driving scenarios consist of video (5 frames), multi-view images (6 views), and single-image, respectively.
Data Preprocess.
Notably, the answers in the raw data consist of more tokens (e.g., detailed driving suggestions), distinguishing them from unimodal knowledge editing tasks, such as those represented by triple-based, which have concise token answers. This introduces the following challenges: 1) Suboptimal Editing Performance. Redundant answers pose difficulties for the implementation of certain editing methods, particularly causal tracing-based approaches such as ROME (Meng et al., 2022). 2) Difficult Evaluation. Evaluating the accuracy of long text sequences often requires external models to score semantic consistency, which not only suffers from limited accuracy but also incurs high computational costs. Consequently, simplifying the answers in the raw data becomes a necessary step, as shown in Figure 3. We prompt Deepseek-v3 (et al, 2024) to condense the original answers (The prompt template is shown in Appendix C). Meanwhile, we retain the original answers as reserved information for potential use in future research.
Reliability Data Construction.
We observe that the data from DriveLM include predictions of vehicle trajectories, which are not feasible for streamlined and generalized data collection. Consequently, we decide to exclude it. For CODA-LM, which contains a significant amount of autonomous driving suggestions and image descriptions, we opt to use Deepseek-v3 to self-generate QA pairs (The prompt template is shown in Appendix C). Subsequently, we shuffle all the data and select a subset to serve as reliability data.
3.2.2 Generality Data Construction
Textual Generality Data.
Given the same driving scenario, the text query is modified to test whether the model truly understands the underlying context. Following prior works, we prompt Deepseek-v3 to rewrite the original query into a semantically similar query while keeping the answer unchanged (See the prompt in Appendix C).
Multimodal Generality Data.
In previous work, text-to-image generation models were employed to create images, or manually curated similar images were used to obtain visually generalized data. However, directly using generative models yields suboptimal performance for the driving domain, which relies exclusively on video data or multiple-view images. The cost of a fully manual selection of similar autonomous driving videos or images is prohibitively high. To address this limitation, we hypothesize that videos corresponding to data with identical questions and answers exhibit similar visual characteristics. Therefore, we only need to match data with identical QA pairs, and then perform non-replacement sampling of two data points per round until exhaustion or only one data point remains. Notably, this process will only occur within the same type of visual data.
3.2.3 Locality Data Construction
To evaluate the capability of editing methods in preserving text locality, we select the Natural Questions dataset (Kwiatkowski et al., 2019), which contains unimodal factual and commonsense knowledge, and randomly sample a subset of it to serve as Textual Locality data. For assessing the preservation of multimodal general capabilities, we select a portion of the VQAv2 dataset (Antol et al., 2015), which includes general visual question answers, to construct Multimodal Locality data.
| Method | Reliability | T-Generality | M-Generality | T-Locality | M-Locality | |
| LLaVA-Onevision | ||||||
| Prompt | 94.25 | 90.18 | 95.04 | 84.86 | 68.13 | |
| AdaLora | 78.01 | 72.76 | 75.84 | 85.51 | 81.12 | |
| GRACE | 100.00 | 28.91 | 28.16 | 100.00 | 100.00 | |
| WISE | 99.10 | 86.97 | 95.78 | 94.18 | 99.98 | |
| Qwen2-VL | ||||||
| Prompt | 90.57 | 84.98 | 90.48 | 89.61 | 72.44 | |
| AdaLora | 79.89 | 75.68 | 78.76 | 82.27 | 69.37 | |
| GRACE | 100.00 | 27.01 | 29.93 | 100.00 | 100.00 | |
| WISE | 94.18 | 85.20 | 91.99 | 94.23 | 99.85 | |
3.3 Quality Control
As outlined in the aforementioned process and illustrated in Figure 3, the entire pipeline is fully automated. However, ensuring data quality through manual verification remains a critical step. To this end, we conduct manual verification and calibration for all processes involving AI models, including answer simplification, QA pairs generation from CODA-LM, textual generality data generation, and scenario categorization. Furthermore, for multimodal generality data, we manually sample 20% of the data to assess similarity and filter out those instances with no resemblance. The data quality control tasks are evenly distributed among three annotators with graduate-level education, who independently performed the calibration. In cases of uncertainty, the annotators first independently judge on whether modifications are necessary, and then resolve any disagreements through a majority vote. The inter-annotator consistency, measured by Fleiss’s Kappa (), demonstrates substantial agreement ( = 0.912), based on a randomly selected sample of 200 annotated data points, indicating a high level of consistency within the range of 0.80 ≤ ≤ 1.00.
4 Evaluation
4.1 Experimental settings
We conduct experiments on the most advanced LMMs to date, which are LLaVA-OneVision Li et al. (2024a) and Qwen2-VL Wang et al. (2024a). Both models utilize Qwen2-7B as their LLM component. Meanwhile, four classical knowledge editing baselines, including Prompt, AdaLora, GRACE, and WISE are employed to update the knowledge of LMMs. Further details can be found in the Appendix B. To comprehensively evaluate the performance of various baselines in ADS-Edit, we test in Single Editing and Lifelong Editing scenarios. Single editing involves updating and evaluating the model immediately after receiving each individual driving data instance. In contrast, lifelong editing represents a more realistic driving scenario, where multiple driving data instances are collected during the vehicle’s moving, requiring continuous integration of this knowledge into the model followed by evaluation. We reasonably assume that the visual processing component (e.g., Vision Transformer) of LMMs provides reliable visual information. Therefore, we focus only on editing the LLM component to modify its understanding of visual input.
4.2 Main Results
4.2.1 Single Editing Results
Memory-based editing methods, such as GRACE and WISE, have proven highly effective in modulating the behavior of LMMs for autonomous driving task prediction. Notably, GRACE achieves a 100% modification rate in both LLaVA-Onevision and Qwen2-VL. Although the ground truth answer is directly provided in the input text, the reliability of the Prompt remains suboptimal. We hypothesize that this is due to the model being affected by interference when processing multimodal driving inputs, which differs from the unimodal scenarios. Under a limited number of training epochs, AdaLora struggles to achieve satisfactory reliability through model parameter updates. Evaluating the generalization of editing methods to other autonomous driving scenarios and queries is essential. However, GRACE exhibits the worst prediction performance on Generality, with accuracy dropping below 30% in both LMMs, compared to its Reliability. This is attributed to the difficulty of GRACE’s codebook in capturing the representational differences of long-sequence multimodal inputs. The remaining three baselines show similar Generality and Reliability results. Among them, WISE and Prompt achieve approximately 90% on Generality, demonstrating their stable generalization capabilities. AdaLora still struggle to generalize the other driving scenarios. Furthermore, M-Generality of the baselines exceeds T-Generality, suggesting that LMMs tend to focus more on text tokens while making less use of visual tokens. This observation may inspire further research on enhancing the efficient utilization of redundant visual representations in LMMs. While preserving the model’s local capabilities, GRACE demonstrates a significant advantage in Locality. WISE, which employs a dual-memory mechanism, also maintains good Locality, achieving nearly 100%. In contrast, AdaLora and Prompt somewhat disrupt the original model’s Locality. Additionally, we observe that both AdaLora and Prompt exhibit lower M-Locality than T-Locality, which is due to the changes in the LMM’s behavior caused by the multimodal inputs in autonomous driving scenarios.
4.2.2 Lifelong Editing Results
In real-world autonomous driving scenarios, sequential knowledge updates are often required. We sequentially test the effects of lifelong editing 1, 250, 500, 750, and 1000 times across four baseline methods: Prompt, AdaLora, GRACE, and WISE. The lifelong editing results of Qwen2-VL and LLaVA-Onevision are shown in Figure 4 and Figure 8, respectively. Regarding Reliability and Generality metrics, the performance of the four baselines shows a gradual decline, albeit with some fluctuations. WISE, as a memory-augmented approach, effectively alleviates the knowledge forgetting phenomenon observed in AdaLora, which relies on parameter updates. Although the Prompt method demonstrates also effective lifelong editing performance at 250 and 500 times updates, it leds to an Out-of-Memory (OOM) error after 750 times updates. Due to the long tokens of multimodal inputs, the codebook struggles to differentiate between distinct representation information, resulting in suboptimal lifelong editing performance for GRACE. Methods that preserve the original parameters, such as WISE and GRACE, achieved 100% performance on the Locality metric. Due to interference from additional inputs, the Prompt method performs poorly on the Locality metric, with even worse results observed in multimodal locality. AdaLoRA employs low-rank matrix-based parameter updates, which minimally affect the original parameters, resulting in a strong performance on the locality metric.
4.3 Analysis
LLaVA-OneVision vs. Qwen2-VL: Which is Easier to Edit?
Although both models use Qwen2 as Large Language Model, LLaVA-Onevision achieve better results across most metrics in terms of reliability and generality. Specifically, on Prompt and WISE methods, Qwen2-VL demonstrates stronger performance in retaining the original predicted answers. Qwen2-VL slightly outperforms LLaVA-OneVision in terms of reliability and generality by using AdaLora. However, locality is significantly compromised compared to LLaVA-OneVision.
Which Scenario Type Knowledges are Easier to Update?
We choose the average generality metric, which is the mean of the text and multimodal generality of Qwen2-VL and LLaVA-OneVision, to assess the performance of the edits. (Since both WISE and GRACE achieve Reliability close to 100%, analyzing the Reliability metric becomes less meaningful. This observation applies to all subsequent analyses in this study.) According to the averaged metrics across three scenarios illustrated in Figure 5, for AdaLora, decision scenario is more challenging to learn compared to perception and understanding scenario, given a fixed number of training epochs. This can be attributed to the higher complexity of decision making data. In contrast, Prompt, GRACE, and WISE show relatively consistent performance metrics across all scenarios, with minimal variation in their generality. Overall, these knowledge editing methods are broadly applicable to various driving scenarios, effectively updating knowledge in LMMs.
Which Data Type is Easier to Edit?
We compute the baseline performance across different types of modality, including video, multi-view images and single image, as shown in Figure 6. WISE demonstrates the highest performance on video data, likely due to its memory mechanism to store and update knowledge in temporal changes in driving scenario. Both Prompt and AdaLoRA exhibit a gradual increase in generality as the length of the visual sequence decreases.
Does Reducing Video Frames Impact the Effectiveness of Knowledge Editing?
To assess the impact of video frames on editing performance, we sequentially test videos with 1 to 5 frames as input in single editing setting. The average generality on video data with 1-4 frames is denoted as the effect of the knowledge editing method under low video frame rate conditions, whereas video data with 5 frames is referred to as the maximum frame rate condition. Fewer video frames correspond to fewer visual tokens, enabling LMMs to process users’ requests more quickly. Although this reduction in frames results in some loss of information, WISE demonstrates even better performance. Furthermore, the results of both Prompt and AdaLora are not significantly affected under conditions of both low and maximum frame rate. This suggests that knowledge editing methods in the autonomous driving domain, particularly for video data, can effectively balance processing speed and performance.
| Low Video Frame rate | Max Video Frame rate | |
| Prompt | 90.14 | 92.61 |
| AdaLora | 70.04 | 74.30 |
| GRACE | 34.33 | 28.54 |
| WISE | 99.49 | 91.37 |
Cases Analysis of Knowledge Editing.
From Figure 7, we observe that the edited LMM tends to maintain better locality in the unimodal domain compared to the multimodal one. Interestingly, despite only editing the language model component of the LMM, the model’s visual understanding is impaired. This underscores the distinct differences between the LMM’s capabilities in visual understanding and text-based knowledge storage.
5 Related Work
5.1 Knowledge Editing
Recent advances in knowledge editing have emerged as a pivotal research direction in knowledge updating (Chen, 2024; Liu et al., 2024b; Wang et al., 2024c; Wu et al., 2024a; Xu et al., 2024; Li and Chu, 2024; Nie et al., 2024; Wei et al., 2025; Zhang et al., 2025; Ali et al., 2025; Hwang et al., 2025; Zhao et al., 2025; Markowitz et al., 2025; Yang et al., 2025; Dong et al., 2025; Pan et al., 2025; Wang et al., 2025; Liu et al., ; Zhang et al., 2024b; Xu et al., 2025; Bi et al., 2025; He et al., 2025; Li et al., 2025; Shen and Huang, 2025). Following Yao et al. (2023), we systematize current methodologies into two principal paradigms: 1) Parameter Modification Approaches: These methods directly alter the model’s internal representations through targeted interventions (Meng et al., 2022, 2023; Fang et al., 2024b; Cai and Cao, 2024; Mitchell et al., 2022; Zhang et al., 2024c; Tan et al., 2024; Hu et al., 2022; Zhang et al., 2023; Dettmers et al., 2023; Feng et al., 2025). 2) Parameter Preservation Approaches: These strategies maintain original parameters while updating new knowledge (Zheng et al., 2023; Jiang et al., 2024; Wang et al., 2024b; Hartvigsen et al., 2023; Li et al., 2025). The emergence of Large Multimodal Models (LMMs) has also advanced the field of multimodal knowledge editing. These works mainly focus on the editing of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024; Pan et al., 2024; Wang et al., 2024d) or multimodal factual knowledge (Li et al., 2024b; Zhang et al., 2024a; Huang et al., 2024a). However, the application of multimodal knowledge editing in the context of domain-specific knowledge, like autonomous driving remains underexplored.
5.2 Large Multi-modal Models for Autonomous Driving
The recent rise of large multimodal models (LMMs) in autonomous driving (Qian et al., 2024; Cui et al., 2024) has spurred the development of advanced frameworks that integrate perception, prediction, and planning. CODA-LM (Li et al., 2024c) and DriveLM (Sima et al., 2024), have attempted to bridge this gap by integrating hierarchical reasoning and graph-structured visual QA frameworks. CODA-LM evaluates vision-language models on corner cases across perception, prediction, and planning stages, while DriveLM explicitly models logical dependencies between driving phases. Meanwhile, LingoQA (Marcu et al., 2024) expands free-form QA capabilities by integrating action justification and scene understanding. However, these models still face challenges in real-world deployment due to static knowledge representations and modality imbalance. To address these issues, knowledge editing (Yao et al., 2023) has emerged as a promising direction, enabling targeted modifications to model behavior without full retraining. Yet, existing datasets and methods fail to account for the unique demands of autonomous driving, such as multimodal coherence and lifelong adaptation to evolving scenarios. This gap motivates our work, ADS-Edit, the first multimodal knowledge editing dataset designed explicitly for autonomous driving systems.
6 Conclusion
In this paper, we present to leverage knowledge editing techniques to address the challenges faced by Large Multimodal Models in autonomous driving scenarios. ADS-Edit, a new benchmark specifically is designed for evaluating knowledge editing methods in this domain. Through extensive experimental analysis, we systematically compare and evaluate the effectiveness of various editing methods, while providing in-depth insights into the underlying reasons for the observed performance variations and failure cases.
Limitations
While this study advances Knowledge Editing for Autonomous Driving System applications, several limitations warrant consideration:
(1) Only VQA Data.
In this work, we only test knowledge editing baselines on visual question answering data and don’t explore other forms of data related to autonomous driving tasks, such as trajectory prediction. How to successfully update predicted location coordinates as knowledge into LMMs and evaluate their generalization will be explored in future work.
(2) Limited Experimental Baselines.
Due to the high cost of LMMs, we did not conduct experiments on larger scale LMMs and discard knowledge editing methods that require more resources, such as MEND and SERAC.
Appendix A Metrics
Reliability.
Reliability is to evaluate the success of behavioral modification in the target driving scenarios. Intuitively, what we need is an updated with . To measure the reliability, we use the editing accuracy, as follows:
where refers to the parameters after editing.
Generality.
Given similar autonomous driving scenarios, generality assesses whether the edited model updates relevant knowledge within to retain the capacity for generalization, so that predict congruent outputs for equivalent inputs (e.g., rephrased textual quries, or similar autonomous driving’s videos). We evaluate text generality and multimodal generality as follows:
where presents the rephrased autonomous driving’s image or videos, refers to the rephrased textual queries, and denotes to in-scope objects of .
Locality.
After updating the model with domain-specific knowledge in autonomous driving, locality is used to evaluate whether knowledge editing preserves the model’s behavior on unrelated knowledge (e.g., unimodal factual knowledge and multimodal commonsense knowledge). Following (Cheng et al., 2023), we employ both text locality and multimodal locality to assess the stability of the editing process:
where and are distinct from . The tuples and represent samples drawn from and , respectively.
Appendix B Baselines
Prompt.
Directly alter the model’s behavior temporarily through prompts.
AdaLora.
LoRA introduces low-rank matrices into Transformer layers, fine-tuning only a small subset of parameters while keeping the majority of the model frozen. Building on LoRA, AdaLoRA (Zhang et al., 2023) improves LoRA by adaptively allocating the parameter budget based on the importance of weight matrices, using SVD to prune unimportant updates. In our work, we employ AdaLoRA as the baseline for efficient knowledge editing.
GRACE.
GRACE (Hartvigsen et al., 2023) introduces a discrete key-value codebook approach, where edits are cached as latent space mappings without altering the original model weights. This method enables thousands of sequential edits by storing corrections in a retrieval-based codebook, ensuring minimal interference with unrelated inputs and strong locality . However, GRACE’s reliance on non-parametric representations limits its ability to generalize beyond memorized edits, as it struggles to integrate new knowledge into the model’s reasoning process.
WISE.
WISE (Wang et al., 2024b) uses a dual-memory architecture comprising a main memory for pretrained knowledge and a side memory for editing. A routing mechanism decides which memory to use during inference. Additionally, knowledge is partitioned into multiple subspaces through sharding to facilitate conflict-free edits, with subsequent merging of the shards into a unified side memory.
Appendix C Benchmark Construction Details
Appendix D Case of ADS-Edit benchmark

### v2

@@
 ADS-Edit: A Multimodal Knowledge Editing Dataset for Autonomous Driving Systems
-Abstract
+Abstract.
 Recent advancements in Large Multimodal Models (LMMs) have shown promise in Autonomous Driving Systems (ADS).
@@
 We conduct comprehensive experiments and derive several interesting conclusions.
-We hope that our work will contribute to the further advancement of knowledge editing applications in the field of autonomous driving111Code and data are available in https://github.com/zjunlp/EasyEdit..
-ADS-Edit: A Multimodal Knowledge Editing Dataset for
-Autonomous Driving Systems
-Chenxi Wang♡11footnotemark: 1, Jizhan Fang♡11footnotemark: 1, Xiang Chen♠††thanks: Equal Contribution., Bozhong Tian♡, Ziwen Xu♡, Huajun Chen♡, Ningyu Zhang♡††thanks: Corresponding author.
-♡Zhejiang University ♠Nanjing University of Aeronautics and Astronautics {sunnywcx,zhangningyu}@zju.edu.cn
-1 Introduction
-The recent Large Multimodal Models (LMMs) (Wang et al., 2024a; Li et al., 2024a; Wu et al., 2024b; Liu et al., 2024c; Team, 2024, 2025) have significantly enhanced capabilities in video understanding and multimodal reasoning.
-As a key foundation, LMMs have also found initial applications in Autonomous Driving Systems (ADS) (Xing et al., 2024; Sima et al., 2024; Huang et al., 2024b; Yao et al., 2024).
+We hope that our work will contribute to the further advancement of knowledge editing applications in the field of autonomous driving111Code and data are available in https://github.com/zjunlp/EasyEdit/blob/main/examples/ADSEdit.md..
+1. Introduction
+The recent Large Multimodal Models (LMMs) (Wang et al., 2024a; Li et al., 2024d; Wu et al., 2024a; Liu et al., 2024c; Team, 2024, 2025) have significantly enhanced capabilities in video understanding and multimodal reasoning.
+As a key foundation, LMMs have also found initial applications in Autonomous Driving Systems (ADS) (Xing et al., 2024; Sima et al., 2024; Huang et al., 2024a; Yao et al., 2024).
 However, the direct application of LMMs in ADS yields suboptimal results.
@@
 These challenges require the model to have the capability of updating knowledge in real-time and continuously.
-To address these challenges, we propose the use of Knowledge Editing, which enables targeted modifications to the model’s behavior without the computational burden of full retraining.
-Unlike traditional fine-tuning, which risks catastrophic forgetting and demands extensive resources, Knowledge Editing facilitate rapid knowledge updates by selectively altering parameters associated with specific factual or contextual knowledge (Yao et al., 2023; Gupta et al., 2024; Zhang et al., 2024d; Fang et al., 2024a; Youssef et al., 2025; Jiang et al., 2025; Liu et al., 2024a; Yao et al., 2025).
-Although Knowledge Editing has been preliminarily explored in the area of LMMs, existing efforts have primarily focused on editing multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024) and multimodal factual knowledge (Zhang et al., 2024a; Huang et al., 2024a).
-To address this limitation, we pioneer the editing of multimodal domain-specific knowledge in the field of autonomous driving.
-Following the principle of leveraging knowledge editing to address the challenges of autonomous driving, we design the ADS-Edit benchmark.
+To address these challenges, we propose the use of Knowledge Editing, which enables targeted modifications to the model’s behavior, eschewing the substantial computational demands associated with full retraining.
+Unlike traditional fine-tuning, which risks catastrophic forgetting and demands extensive resources, Knowledge Editing facilitate rapid knowledge updates by selectively altering parameters associated with specific factual or contextual knowledge (Yao et al., 2023; Gupta et al., 2024; Zhang et al., 2024b; Fang et al., 2024a; Youssef et al., 2025; Jiang et al., 2025; Liu et al., 2024a; Yao et al., 2025).
+While preliminary explorations of Knowledge Editing in the area of LMMs have predominantly concentrated on the manipulation of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024) and multimodal factual knowledge (Zhang et al., 2024d; Huang et al., 2024b), we pioneer the editing of multimodal domain-specific knowledge in the field of autonomous driving.
+Following the principle of leveraging knowledge editing to address the challenges of autonomous driving system, we propose the ADS-Edit benchmark.
 This benchmark encompasses three real-world scenarios: perception, understanding, and decision making.
@@
 Furthermore, we establish comprehensive evaluation metrics for knowledge editing in autonomous driving scenarios.
-We evaluate four commonly used Knowledge Editing baselines under both single editing and lifelong editing scenarios, such as Prompt, AdaLora (Zhang et al., 2023), GRACE (Hartvigsen et al., 2023), and WISE (Wang et al., 2024b).
-Thirdly, through our analysis, we obtained a series of interesting findings, including the universality of knowledge editing methods in updating knowledge across various scenarios, their ability to balance editing effectiveness and processing speed, particularly in the context of video data, and the remaining limitations in locality.
+We evaluate four commonly used Knowledge Editing baselines under both single editing and lifelong editing scenarios, such as Prompt, AdaLora (Zhang et al., 2023), GRACE (Hartvigsen et al., 2023), and WISE (Wang et al., 2024d).
+Through our analysis, we obtained a series of interesting findings, including the universality of knowledge editing methods in updating knowledge across various scenarios, their ability to balance editing effectiveness and processing speed, particularly in the context of video data, and the remaining limitations in locality.
+Moreover, in real-world evaluation scenarios, it is evident that current editing techniques require further refinement and enhancement to meet the demands of autonomous driving systems.
 In general, we conclude our contributions as:
@@
 We test four Knowledge Editing baselines under both single editing and lifelong editing settings and analyze several interesting results.
-2 Background
+2. Background
 LMMs have been applied in autonomous driving, typically after being trained on carefully curated driving datasets.
@@
 Overall, there is a critical need for a framework that enables LMMs to rapidly and sequentially update knowledges when applied to autonomous driving.
-We propose leveraging knowledge editing techniques to address challenges in autonomous driving scenarios, defined as follows: Let be a LMM and be an edited model.
-Given the user’s inputs (Includes text and autonomous driving’s multimodal inputs , such as image or videos) and the editing target , the edited model is expected to modify LMM’s outputs within the editing scope to match the editing target, while preserving the original model’s output for the inputs outside the editing scope.
-The specific definition is as follows:
-where and means the in-scope of the editing inputs.
-3 Bechmark Construction
-3.1 Desgin Principle
+We propose leveraging knowledge editing techniques to address the challenges encountered in autonomous driving scenarios.
+Formally, let denote a multimodal large model (LMM), and let represent the edited model.
+Given a user input , consisting of textual input and multimodal signals from autonomous driving (e.g., images or videos), along with a desired editing target , the goal is to update the model such that .
+The edited model should modify the output of the original LMM within the intended editing scope while preserving its original behavior on inputs outside that scope.
+The formal definition is as follows:
+where and means in-scope of editing inputs.
+3. Bechmark Construction
+3.1. Desgin Principle
 To construct a comprehensive benchmark, we propose a tri-axis design principle.
@@
 The statistics of scenario types is shown in Figure 2.
+Video | Multi-view | Single | All | |
+Train | 1,926 | 960 | 1,093 | 3,979 |
+Test | 481 | 239 | 358 | 1,078 |
 Data Type.
@@
 The statistics of different types of data is shown in Table 1.
-| Video | Multi-view | Single | All | |
-| Train | 1,926 | 960 | 1,093 | 3,979 |
-| Test | 481 | 239 | 358 | 1,078 |
 Metrics.
@@
 Locality measures whether methods alter unrelated knowledge after updating with domain-specific knowledge of autonomous driving.
-3.2 Data Collection
+3.2. Data Collection
 Inspired by knowledge editing applications in unimodal settings, the constructed ADS-Edit dataset is composed of visual question answering data.
 The construction process of the ADS dataset is illustrated in Figure 3.
-3.2.1 Reliability Data Construction
-We select three well-known datasets of the autonomous driving system: LingoQA (Marcu et al., 2024), DriveLM dataset (Sima et al., 2024), and CODA-LM dataset (Li et al., 2024d) as raw data.
-Their autonomous driving scenarios consist of video (5 frames), multi-view images (6 views), and single-image, respectively.
-Data Preprocess.
-Notably, the answers in the raw data consist of more tokens (e.g., detailed driving suggestions), distinguishing them from unimodal knowledge editing tasks, such as those represented by triple-based, which have concise token answers.
+3.2.1. Data Preprocess.
+We select three well-known datasets of the autonomous driving system: LingoQA (Marcu et al., 2024), DriveLM dataset (Sima et al., 2024), and CODA-LM dataset (Li et al., 2024b) as raw data.
+These datasets encompass a variety of autonomous driving scenarios, including video (5 frames), multi-view images (6 perspectives), and single images, respectively.
+Notably, the answers in the raw data consist of more tokens (e.g., detailed driving suggestions), distinguishing them from unimodal knowledge editing tasks, such as those represented by triple-based formats, which typically feature concise token responses.
 This introduces the following challenges: 1) Suboptimal Editing Performance.
@@
 Consequently, simplifying the answers in the raw data becomes a necessary step, as shown in Figure 3.
-We prompt Deepseek-v3 (et al, 2024) to condense the original answers (The prompt template is shown in Appendix C).
-Meanwhile, we retain the original answers as reserved information for potential use in future research.
-Reliability Data Construction.
+We prompt Deepseek-v3 (et al, 2024) to condense the original answers.Meanwhile, we retain the original answers as meta information for potential use in future research.
+3.2.2. Reliability Data Construction
 We observe that the data from DriveLM include predictions of vehicle trajectories, which are not feasible for streamlined and generalized data collection.
 Consequently, we decide to exclude it.
-For CODA-LM, which contains a significant amount of autonomous driving suggestions and image descriptions, we opt to use Deepseek-v3 to self-generate QA pairs (The prompt template is shown in Appendix C).
-Subsequently, we shuffle all the data and select a subset to serve as reliability data.
-3.2.2 Generality Data Construction
+For CODA-LM, which contains a significant amount of autonomous driving suggestions and image descriptions, we opt to use Deepseek-v3 to self-generate QA pairs.We then shuffle all the data and select a subset to serve as reliability data.
+Furthermore, to present the metadata from multiple perspectives, we categorize the selected data by scenario types, comprising three main categories and eleven subcategories.
+Specifically, we prompt DeepSeek-v3 to perform classification, which is then verified through rigorous manual inspection, as detailed in Section 3.3.
+Method | Reliability | T-Generality | M-Generality | T-Locality | M-Locality | ||||||
+Edit. | Real. | Edit. | Real. | Edit. | Real. | Edit. | Real. | Edit. | Real. | ||
+LLaVA-Onevision | |||||||||||
+Prompt | 94.25 | 97.68 | 90.18 | 94.71 | 95.04 | 97.21 | 84.47 | 11.87 | 80.35 | 79.13 | |
+AdaLora | 78.01 | 56.68 | 72.76 | 54.36 | 75.84 | 53.34 | 85.51 | 40.54 | 81.12 | 69.67 | |
+GRACE | 100.00 | 89.52 | 28.91 | 45.83 | 28.16 | 45.45 | 100.00 | 100.00 | 100.00 | 100.00 | |
+WISE | 99.10 | 75.97 | 86.97 | 70.87 | 95.78 | 73.56 | 94.18 | 56.96 | 99.98 | 82.84 | |
+Qwen2-VL | |||||||||||
+Prompt | 90.57 | 97.77 | 84.98 | 94.53 | 90.48 | 97.03 | 89.61 | 61.04 | 72.44 | 66.14 | |
+AdaLora | 79.89 | 65.03 | 75.68 | 61.69 | 78.76 | 63.64 | 82.27 | 41.37 | 69.37 | 57.33 | |
+GRACE | 100.00 | 51.67 | 27.01 | 47.96 | 29.93 | 52.41 | 100.00 | 100.00 | 100.00 | 100.00 | |
+WISE | 94.18 | 62.98 | 85.20 | 57.14 | 91.99 | 62.15 | 94.23 | 70.13 | 99.85 | 71.33 |
+3.2.3. Generality Data Construction
 Textual Generality Data.
 Given the same driving scenario, the text query is modified to test whether the model truly understands the underlying context.
-Following prior works, we prompt Deepseek-v3 to rewrite the original query into a semantically similar query while keeping the answer unchanged (See the prompt in Appendix C).
+Following prior works, we prompt Deepseek-v3 to rewrite the original query into a semantically similar query while keeping the answer unchanged.
 Multimodal Generality Data.
@@
 Notably, this process will only occur within the same type of visual data.
-3.2.3 Locality Data Construction
-To evaluate the capability of editing methods in preserving text locality, we select the Natural Questions dataset (Kwiatkowski et al., 2019), which contains unimodal factual and commonsense knowledge, and randomly sample a subset of it to serve as Textual Locality data.
-For assessing the preservation of multimodal general capabilities, we select a portion of the VQAv2 dataset (Antol et al., 2015), which includes general visual question answers, to construct Multimodal Locality data.
-| Method | Reliability | T-Generality | M-Generality | T-Locality | M-Locality | |
-| LLaVA-Onevision | ||||||
-| Prompt | 94.25 | 90.18 | 95.04 | 84.86 | 68.13 | |
-| AdaLora | 78.01 | 72.76 | 75.84 | 85.51 | 81.12 | |
-| GRACE | 100.00 | 28.91 | 28.16 | 100.00 | 100.00 | |
-| WISE | 99.10 | 86.97 | 95.78 | 94.18 | 99.98 | |
-| Qwen2-VL | ||||||
-| Prompt | 90.57 | 84.98 | 90.48 | 89.61 | 72.44 | |
-| AdaLora | 79.89 | 75.68 | 78.76 | 82.27 | 69.37 | |
-| GRACE | 100.00 | 27.01 | 29.93 | 100.00 | 100.00 | |
-| WISE | 94.18 | 85.20 | 91.99 | 94.23 | 99.85 | |
-3.3 Quality Control
+3.2.4. Locality Data Construction
+To evaluate the ability of editing methods to preserve textual locality, we utilize the Natural Questions dataset (Kwiatkowski et al., 2019), which encompasses unimodal factual and commonsense knowledge.
+A random subset of this dataset is sampled to constitute the Textual Locality data.
+For assessing the preservation of general multimodal capabilities, we select a portion of the VQAv2 dataset (Antol et al., 2015), comprising general visual question-answer pairs, to form the Multimodal Locality data.
+3.3. Quality Control
 As outlined in the aforementioned process and illustrated in Figure 3, the entire pipeline is fully automated.
@@
 Furthermore, for multimodal generality data, we manually sample 20% of the data to assess similarity and filter out those instances with no resemblance.
+To ensure the consistency between the rationales in the generality data answers and the reliability data, we filter out mismatched samples by referring to the original answers.
 The data quality control tasks are evenly distributed among three annotators with graduate-level education, who independently performed the calibration.
 In cases of uncertainty, the annotators first independently judge on whether modifications are necessary, and then resolve any disagreements through a majority vote.
-The inter-annotator consistency, measured by Fleiss’s Kappa (), demonstrates substantial agreement ( = 0.912), based on a randomly selected sample of 200 annotated data points, indicating a high level of consistency within the range of 0.80 ≤ ≤ 1.00.
-4 Evaluation
-4.1 Experimental settings
-We conduct experiments on the most advanced LMMs to date, which are LLaVA-OneVision Li et al.
-(2024a) and Qwen2-VL Wang et al.
-(2024a).
+The inter-annotator consistency, measured by Fleiss’s Kappa (), demonstrates substantial agreement ( = 0.802), based on a randomly selected sample of 200 annotated data points, indicating a high level of consistency within the range of 0.80 ≤ ≤ 1.00.
+4. Evaluation
+4.1. Experimental settings
+We conduct experiments on the most advanced LMMs to date, which are LLaVA-OneVision (Li et al., 2024d) and Qwen2-VL (Wang et al., 2024a).
 Both models utilize Qwen2-7B as their LLM component.
 Meanwhile, four classical knowledge editing baselines, including Prompt, AdaLora, GRACE, and WISE are employed to update the knowledge of LMMs.
-Further details can be found in the Appendix B.
-To comprehensively evaluate the performance of various baselines in ADS-Edit, we test in Single Editing and Lifelong Editing scenarios.
+Our analysis covers both the Single Editing and Lifelong Editing settings to thoroughly assess the performance of different baselines in ADS-Edit.
 Single editing involves updating and evaluating the model immediately after receiving each individual driving data instance.
@@
 Therefore, we focus only on editing the LLM component to modify its understanding of visual input.
-4.2 Main Results
-4.2.1 Single Editing Results
+To comprehensively evaluate the effectiveness of editing methods on ADS, we adopt two evaluation metrics: (1) Editing evaluation (Edit), where tokens are generated in a teacher-forcing manner and compared at the token level (Meng et al., 2022).
+(2) Real-world evaluation (Real), where the edited model performs free-form generation, and a LLM is prompted to assess the semantic consistency between the generated outputs and the ground-truth answers (Yang et al., 2025a).
+Due to cost and resource constraints, we report the Real metric only for the main single-edit experiment, while the remaining experiments are evaluated using the Edit metric by default.
+4.2. Main Results
+We present the results for the single-edit scenario below; the results for lifelong editing can be found on Github222https://github.com/zjunlp/EasyEdit/blob/main/examples/ADSEdit.md.
+Results of Editing evaluation.
 Memory-based editing methods, such as GRACE and WISE, have proven highly effective in modulating the behavior of LMMs for autonomous driving task prediction.
-Notably, GRACE achieves a 100% modification rate in both LLaVA-Onevision and Qwen2-VL.
-Although the ground truth answer is directly provided in the input text, the reliability of the Prompt remains suboptimal.
-We hypothesize that this is due to the model being affected by interference when processing multimodal driving inputs, which differs from the unimodal scenarios.
-Under a limited number of training epochs, AdaLora struggles to achieve satisfactory reliability through model parameter updates.
-Evaluating the generalization of editing methods to other autonomous driving scenarios and queries is essential.
-However, GRACE exhibits the worst prediction performance on Generality, with accuracy dropping below 30% in both LMMs, compared to its Reliability.
-This is attributed to the difficulty of GRACE’s codebook in capturing the representational differences of long-sequence multimodal inputs.
-The remaining three baselines show similar Generality and Reliability results.
-Among them, WISE and Prompt achieve approximately 90% on Generality, demonstrating their stable generalization capabilities.
-AdaLora still struggle to generalize the other driving scenarios.
-Furthermore, M-Generality of the baselines exceeds T-Generality, suggesting that LMMs tend to focus more on text tokens while making less use of visual tokens.
-This observation may inspire further research on enhancing the efficient utilization of redundant visual representations in LMMs.
-While preserving the model’s local capabilities, GRACE demonstrates a significant advantage in Locality.
-WISE, which employs a dual-memory mechanism, also maintains good Locality, achieving nearly 100%.
-In contrast, AdaLora and Prompt somewhat disrupt the original model’s Locality.
-Additionally, we observe that both AdaLora and Prompt exhibit lower M-Locality than T-Locality, which is due to the changes in the LMM’s behavior caused by the multimodal inputs in autonomous driving scenarios.
-4.2.2 Lifelong Editing Results
-In real-world autonomous driving scenarios, sequential knowledge updates are often required.
-We sequentially test the effects of lifelong editing 1, 250, 500, 750, and 1000 times across four baseline methods: Prompt, AdaLora, GRACE, and WISE.
-The lifelong editing results of Qwen2-VL and LLaVA-Onevision are shown in Figure 4 and Figure 8, respectively.
-Regarding Reliability and Generality metrics, the performance of the four baselines shows a gradual decline, albeit with some fluctuations.
-WISE, as a memory-augmented approach, effectively alleviates the knowledge forgetting phenomenon observed in AdaLora, which relies on parameter updates.
-Although the Prompt method demonstrates also effective lifelong editing performance at 250 and 500 times updates, it leds to an Out-of-Memory (OOM) error after 750 times updates.
-Due to the long tokens of multimodal inputs, the codebook struggles to differentiate between distinct representation information, resulting in suboptimal lifelong editing performance for GRACE.
-Methods that preserve the original parameters, such as WISE and GRACE, achieved 100% performance on the Locality metric.
-Due to interference from additional inputs, the Prompt method performs poorly on the Locality metric, with even worse results observed in multimodal locality.
-AdaLoRA employs low-rank matrix-based parameter updates, which minimally affect the original parameters, resulting in a strong performance on the locality metric.
-4.3 Analysis
+Notably, GRACE achieves a 100% modification rate in both models, yet performs poorly in terms of generality.
+WISE demonstrates a well-balanced performance across reliability, generality, and locality.
+Although the ground truth answer is directly provided in the input text, the reliability and generality of the Prompt remains suboptimal.
+Meanwhile, Prompt performs poorly on the locality, which can be attributed to its sensitivity to the ground-truth answers in the input.
+Under a limited resources, AdaLora struggles to achieve satisfactory reliability and generality through model parameter updates.
+Moreover, parameter modifications significantly impact Adalora’s performance on the locality metric.
+Results of Real-world evaluation.
+Although Prompt underperforms GRACE and WISE on conventional editing metrics, it achieves the best performance in terms of reliability and generalization under real-world evaluation settings.
+This is primarily because the model does not strictly replicate the ground-truth answers but instead generates outputs that are semantically aligned with them.
+In contrast, the other three editing methods exhibit a noticeable decline in performance when transitioning from editing metrics to real-world evaluation, indicating limitations in their reliability and generalization for ADS applications.
+In terms of locality metrics, GRACE and WISE demonstrate strong performance, whereas Prompt underperforms.
+4.3. Analysis
 LLaVA-OneVision vs. Qwen2-VL: Which is Easier to Edit?
@@
 However, locality is significantly compromised compared to LLaVA-OneVision.
+Furthermore, in real-world evaluation scores, we observe that Qwen2-VL tends to retain its original outputs and exhibits resistance to editing.
 Which Scenario Type Knowledges are Easier to Update?
@@
 Overall, these knowledge editing methods are broadly applicable to various driving scenarios, effectively updating knowledge in LMMs.
+Low Video Frame rate | Max Video Frame rate | |
+Prompt | 90.14 | 92.61 |
+AdaLora | 70.04 | 74.30 |
+GRACE | 34.33 | 28.54 |
+WISE | 99.49 | 91.37 |
 Which Data Type is Easier to Edit?
@@
 This suggests that knowledge editing methods in the autonomous driving domain, particularly for video data, can effectively balance processing speed and performance.
-| Low Video Frame rate | Max Video Frame rate | |
-| Prompt | 90.14 | 92.61 |
-| AdaLora | 70.04 | 74.30 |
-| GRACE | 34.33 | 28.54 |
-| WISE | 99.49 | 91.37 |
 Cases Analysis of Knowledge Editing.
@@
 This underscores the distinct differences between the LMM’s capabilities in visual understanding and text-based knowledge storage.
-5 Related Work
-5.1 Knowledge Editing
-Recent advances in knowledge editing have emerged as a pivotal research direction in knowledge updating (Chen, 2024; Liu et al., 2024b; Wang et al., 2024c; Wu et al., 2024a; Xu et al., 2024; Li and Chu, 2024; Nie et al., 2024; Wei et al., 2025; Zhang et al., 2025; Ali et al., 2025; Hwang et al., 2025; Zhao et al., 2025; Markowitz et al., 2025; Yang et al., 2025; Dong et al., 2025; Pan et al., 2025; Wang et al., 2025; Liu et al., ; Zhang et al., 2024b; Xu et al., 2025; Bi et al., 2025; He et al., 2025; Li et al., 2025; Shen and Huang, 2025).
+5. Related Work
+5.1. Knowledge Editing
+Recent advances in knowledge editing have emerged as a pivotal research direction in knowledge updating (Chen, 2024; Liu et al., 2024b; Wang et al., 2024b; Wu et al., 2024b; Xu et al., 2024; Li and Chu, 2024; Nie et al., 2024; Wei et al., 2025; Zhang et al., 2025; Ali et al., 2025; Hwang et al., 2025; Zhao et al., 2025; Markowitz et al., 2025; Yang et al., 2025b; Dong et al., 2025; Pan et al., 2025; Wang et al., 2025; Liu et al., [n.
+d.]; Zhang et al., 2024c; Xu et al., 2025; Bi et al., 2025; He et al., 2025; Li et al., 2025; Shen and Huang, 2025; Yang et al., 2025a).
 Following Yao et al.
-(2023), we systematize current methodologies into two principal paradigms: 1) Parameter Modification Approaches: These methods directly alter the model’s internal representations through targeted interventions (Meng et al., 2022, 2023; Fang et al., 2024b; Cai and Cao, 2024; Mitchell et al., 2022; Zhang et al., 2024c; Tan et al., 2024; Hu et al., 2022; Zhang et al., 2023; Dettmers et al., 2023; Feng et al., 2025).
-2) Parameter Preservation Approaches: These strategies maintain original parameters while updating new knowledge (Zheng et al., 2023; Jiang et al., 2024; Wang et al., 2024b; Hartvigsen et al., 2023; Li et al., 2025).
+(2023), we systematize current methodologies into two principal paradigms: (1) Parameter Modification Approaches: These methods directly alter the model’s internal representations through targeted interventions (Meng et al., 2022, 2023; Fang et al., 2024b; Cai and Cao, 2024; Mitchell et al., 2022; Zhang et al., 2024a; Tan et al., 2024; Hu et al., 2022; Zhang et al., 2023; Dettmers et al., 2023; Feng et al., 2025).
+(2) Parameter Preservation Approaches: These strategies maintain original parameters while updating new knowledge (Zheng et al., 2023; Jiang et al., 2024; Wang et al., 2024d; Hartvigsen et al., 2023; Li et al., 2025).
 The emergence of Large Multimodal Models (LMMs) has also advanced the field of multimodal knowledge editing.
-These works mainly focus on the editing of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024; Pan et al., 2024; Wang et al., 2024d) or multimodal factual knowledge (Li et al., 2024b; Zhang et al., 2024a; Huang et al., 2024a).
+These works mainly focus on the editing of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024; Pan et al., 2024; Wang et al., 2024c) or multimodal factual knowledge (Li et al., 2024a; Zhang et al., 2024d; Huang et al., 2024b).
 However, the application of multimodal knowledge editing in the context of domain-specific knowledge, like autonomous driving remains underexplored.
-5.2 Large Multi-modal Models for Autonomous Driving
-The recent rise of large multimodal models (LMMs) in autonomous driving (Qian et al., 2024; Cui et al., 2024) has spurred the development of advanced frameworks that integrate perception, prediction, and planning.
-CODA-LM (Li et al., 2024c) and DriveLM (Sima et al., 2024), have attempted to bridge this gap by integrating hierarchical reasoning and graph-structured visual QA frameworks.
-CODA-LM evaluates vision-language models on corner cases across perception, prediction, and planning stages, while DriveLM explicitly models logical dependencies between driving phases.
+5.2. Large Multi-modal Models for Autonomous Driving
+The recent rise of large multimodal models (LMMs) in autonomous driving (Qian et al., 2024; Cui et al., 2024) has spurred the development of advanced frameworks.
+CODA-LM (Li et al., 2024c) and DriveLM (Sima et al., 2024) integrate hierarchical reasoning and graph-structured visual QA frameworks.
 Meanwhile, LingoQA (Marcu et al., 2024) expands free-form QA capabilities by integrating action justification and scene understanding.
 However, these models still face challenges in real-world deployment due to static knowledge representations and modality imbalance.
-To address these issues, knowledge editing (Yao et al., 2023) has emerged as a promising direction, enabling targeted modifications to model behavior without full retraining.
-Yet, existing datasets and methods fail to account for the unique demands of autonomous driving, such as multimodal coherence and lifelong adaptation to evolving scenarios.
-This gap motivates our work, ADS-Edit, the first multimodal knowledge editing dataset designed explicitly for autonomous driving systems.
-6 Conclusion
-In this paper, we present to leverage knowledge editing techniques to address the challenges faced by Large Multimodal Models in autonomous driving scenarios.
-ADS-Edit, a new benchmark specifically is designed for evaluating knowledge editing methods in this domain.
-Through extensive experimental analysis, we systematically compare and evaluate the effectiveness of various editing methods, while providing in-depth insights into the underlying reasons for the observed performance variations and failure cases.
-Limitations
-While this study advances Knowledge Editing for Autonomous Driving System applications, several limitations warrant consideration:
-(1) Only VQA Data.
-In this work, we only test knowledge editing baselines on visual question answering data and don’t explore other forms of data related to autonomous driving tasks, such as trajectory prediction.
-How to successfully update predicted location coordinates as knowledge into LMMs and evaluate their generalization will be explored in future work.
-(2) Limited Experimental Baselines.
-Due to the high cost of LMMs, we did not conduct experiments on larger scale LMMs and discard knowledge editing methods that require more resources, such as MEND and SERAC.
+6. Conclusion
+In this paper, we propose leveraging knowledge editing techniques to address the challenges faced by Large Multimodal Models in autonomous driving scenarios.
+To this end, we introduce ADS-Edit, a novel benchmark specifically designed to evaluate knowledge editing methods within this domain.
+Through extensive experimental analysis, we systematically compare and assess the effectiveness of various editing approaches, providing comprehensive insights into the underlying causes of performance variations and failure cases.
+Moreover, our findings highlight that developing editing methods that are both well-balanced and high-performing remains a significant challenge in the autonomous driving systems.
 Appendix A Metrics
@@
 WISE.
-WISE (Wang et al., 2024b) uses a dual-memory architecture comprising a main memory for pretrained knowledge and a side memory for editing.
+WISE (Wang et al., 2024d) uses a dual-memory architecture comprising a main memory for pretrained knowledge and a side memory for editing.
 A routing mechanism decides which memory to use during inference.
 Additionally, knowledge is partitioned into multiple subspaces through sharding to facilitate conflict-free edits, with subsequent merging of the shards into a unified side memory.
-Appendix C Benchmark Construction Details
-Appendix D Case of ADS-Edit benchmark
+Appendix C Lifelong Editing Results
+In real-world autonomous driving scenarios, sequential knowledge updates are often required.
+We sequentially test the effects of lifelong editing 1, 250, 500, 750, and 1000 times across four baseline methods: Prompt, AdaLora, GRACE, and WISE.
+The lifelong editing results of Qwen2-VL and LLaVA-Onevision are shown in Figure 7 and Figure 8, respectively.
+Regarding Reliability and Generality metrics, the performance of the four baselines shows a gradual decline, albeit with some fluctuations.
+WISE, as a memory-augmented approach, effectively alleviates the knowledge forgetting phenomenon observed in AdaLora, which relies on parameter updates.
+Although the Prompt method demonstrates also effective lifelong editing performance at 250 and 500 times updates, it leds to an Out-of-Memory (OOM) error after 750 times updates.
+Due to the long tokens of multimodal inputs, the codebook struggles to differentiate between distinct representation information, resulting in suboptimal lifelong editing performance for GRACE.
+Methods that preserve the original parameters, such as WISE and GRACE, achieved 100% performance on the Locality metric.
+Due to interference from additional inputs, the Prompt method performs poorly on the Locality metric, with even worse results observed in multimodal locality.
+AdaLoRA employs low-rank matrix-based parameter updates, which minimally affect the original parameters, resulting in a strong performance on the locality metric.
+Appendix D Benchmark Construction Details
+Appendix E Case of ADS-Edit benchmark

### v3

@@
 1. Introduction
-The recent Large Multimodal Models (LMMs) (Wang et al., 2024a; Li et al., 2024d; Wu et al., 2024a; Liu et al., 2024c; Team, 2024, 2025) have significantly enhanced capabilities in video understanding and multimodal reasoning.
+The recent Large Multimodal Models (LMMs) (Wang et al., 2024a; Li et al., 2024d; Team, 2025) have significantly enhanced capabilities in video understanding and multimodal reasoning.
 As a key foundation, LMMs have also found initial applications in Autonomous Driving Systems (ADS) (Xing et al., 2024; Sima et al., 2024; Huang et al., 2024a; Yao et al., 2024).
@@
 Unlike traditional fine-tuning, which risks catastrophic forgetting and demands extensive resources, Knowledge Editing facilitate rapid knowledge updates by selectively altering parameters associated with specific factual or contextual knowledge (Yao et al., 2023; Gupta et al., 2024; Zhang et al., 2024b; Fang et al., 2024a; Youssef et al., 2025; Jiang et al., 2025; Liu et al., 2024a; Yao et al., 2025).
-While preliminary explorations of Knowledge Editing in the area of LMMs have predominantly concentrated on the manipulation of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024) and multimodal factual knowledge (Zhang et al., 2024d; Huang et al., 2024b), we pioneer the editing of multimodal domain-specific knowledge in the field of autonomous driving.
+While preliminary explorations of Knowledge Editing in the area of LMMs have predominantly concentrated on the manipulation of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024) and multimodal factual knowledge (Zhang et al., 2024c; Huang et al., 2024b), we pioneer the editing of multimodal domain-specific knowledge in the field of autonomous driving.
 Following the principle of leveraging knowledge editing to address the challenges of autonomous driving system, we propose the ADS-Edit benchmark.
@@
 Furthermore, we establish comprehensive evaluation metrics for knowledge editing in autonomous driving scenarios.
-We evaluate four commonly used Knowledge Editing baselines under both single editing and lifelong editing scenarios, such as Prompt, AdaLora (Zhang et al., 2023), GRACE (Hartvigsen et al., 2023), and WISE (Wang et al., 2024d).
+We evaluate four commonly used Knowledge Editing baselines under both single editing and lifelong editing scenarios, such as Prompt, AdaLora (Zhang et al., 2023), GRACE (Hartvigsen et al., 2023), and WISE (Wang et al., 2024c).
 Through our analysis, we obtained a series of interesting findings, including the universality of knowledge editing methods in updating knowledge across various scenarios, their ability to balance editing effectiveness and processing speed, particularly in the context of video data, and the remaining limitations in locality.
@@
 To comprehensively evaluate the effectiveness of editing methods on ADS, we adopt two evaluation metrics: (1) Editing evaluation (Edit), where tokens are generated in a teacher-forcing manner and compared at the token level (Meng et al., 2022).
-(2) Real-world evaluation (Real), where the edited model performs free-form generation, and a LLM is prompted to assess the semantic consistency between the generated outputs and the ground-truth answers (Yang et al., 2025a).
+(2) Real-world evaluation (Real), where the edited model performs free-form generation, and a LLM is prompted to assess the semantic consistency between the generated outputs and the ground-truth answers (Yang et al., 2025).
 Due to cost and resource constraints, we report the Real metric only for the main single-edit experiment, while the remaining experiments are evaluated using the Edit metric by default.
@@
 5.1. Knowledge Editing
-Recent advances in knowledge editing have emerged as a pivotal research direction in knowledge updating (Chen, 2024; Liu et al., 2024b; Wang et al., 2024b; Wu et al., 2024b; Xu et al., 2024; Li and Chu, 2024; Nie et al., 2024; Wei et al., 2025; Zhang et al., 2025; Ali et al., 2025; Hwang et al., 2025; Zhao et al., 2025; Markowitz et al., 2025; Yang et al., 2025b; Dong et al., 2025; Pan et al., 2025; Wang et al., 2025; Liu et al., [n.
-d.]; Zhang et al., 2024c; Xu et al., 2025; Bi et al., 2025; He et al., 2025; Li et al., 2025; Shen and Huang, 2025; Yang et al., 2025a).
+Recent advances in knowledge editing have emerged as a pivotal research direction in knowledge updating (Chen, 2024; Liu et al., 2024b; Wang et al., 2024b; Wu et al., 2024; Xu et al., 2024).
 Following Yao et al.
-(2023), we systematize current methodologies into two principal paradigms: (1) Parameter Modification Approaches: These methods directly alter the model’s internal representations through targeted interventions (Meng et al., 2022, 2023; Fang et al., 2024b; Cai and Cao, 2024; Mitchell et al., 2022; Zhang et al., 2024a; Tan et al., 2024; Hu et al., 2022; Zhang et al., 2023; Dettmers et al., 2023; Feng et al., 2025).
-(2) Parameter Preservation Approaches: These strategies maintain original parameters while updating new knowledge (Zheng et al., 2023; Jiang et al., 2024; Wang et al., 2024d; Hartvigsen et al., 2023; Li et al., 2025).
+(2023), we systematize current methodologies into two principal paradigms: (1) Parameter Modification Approaches: These methods directly alter the model’s internal representations through targeted interventions (Meng et al., 2022, 2023; Fang et al., 2024b; Cai and Cao, 2024; Mitchell et al., 2022; Zhang et al., 2024a; Hu et al., 2022; Zhang et al., 2023; Dettmers et al., 2023).
+(2) Parameter Preservation Approaches: These strategies maintain original parameters while updating new knowledge (Zheng et al., 2023; Wang et al., 2024c; Hartvigsen et al., 2023).
 The emergence of Large Multimodal Models (LMMs) has also advanced the field of multimodal knowledge editing.
-These works mainly focus on the editing of multimodal common knowledge (Cheng et al., 2023; Ma et al., 2024; Pan et al., 2024; Wang et al., 2024c) or multimodal factual knowledge (Li et al., 2024a; Zhang et al., 2024d; Huang et al., 2024b).
+These works mainly focus on the editing of multimodal common knowledge (Cheng et al., 2023) or multimodal factual knowledge (Li et al., 2024a; Huang et al., 2024b).
 However, the application of multimodal knowledge editing in the context of domain-specific knowledge, like autonomous driving remains underexplored.
 5.2. Large Multi-modal Models for Autonomous Driving
-The recent rise of large multimodal models (LMMs) in autonomous driving (Qian et al., 2024; Cui et al., 2024) has spurred the development of advanced frameworks.
+The recent rise of large multimodal models (LMMs) in autonomous driving (Qian et al., 2024) has spurred the development of advanced frameworks.
 CODA-LM (Li et al., 2024c) and DriveLM (Sima et al., 2024) integrate hierarchical reasoning and graph-structured visual QA frameworks.
@@
 Moreover, our findings highlight that developing editing methods that are both well-balanced and high-performing remains a significant challenge in the autonomous driving systems.
-Appendix A Metrics
-Reliability.
-Reliability is to evaluate the success of behavioral modification in the target driving scenarios.
-Intuitively, what we need is an updated with .
-To measure the reliability, we use the editing accuracy, as follows:
-where refers to the parameters after editing.
-Generality.
-Given similar autonomous driving scenarios, generality assesses whether the edited model updates relevant knowledge within to retain the capacity for generalization, so that predict congruent outputs for equivalent inputs (e.g., rephrased textual quries, or similar autonomous driving’s videos).
-We evaluate text generality and multimodal generality as follows:
-where presents the rephrased autonomous driving’s image or videos, refers to the rephrased textual queries, and denotes to in-scope objects of .
-Locality.
-After updating the model with domain-specific knowledge in autonomous driving, locality is used to evaluate whether knowledge editing preserves the model’s behavior on unrelated knowledge (e.g., unimodal factual knowledge and multimodal commonsense knowledge).
-Following (Cheng et al., 2023), we employ both text locality and multimodal locality to assess the stability of the editing process:
-where and are distinct from . The tuples and represent samples drawn from and , respectively.
-Appendix B Baselines
-Prompt.
-Directly alter the model’s behavior temporarily through prompts.
-AdaLora.
-LoRA introduces low-rank matrices into Transformer layers, fine-tuning only a small subset of parameters while keeping the majority of the model frozen.
-Building on LoRA, AdaLoRA (Zhang et al., 2023) improves LoRA by adaptively allocating the parameter budget based on the importance of weight matrices, using SVD to prune unimportant updates.
-In our work, we employ AdaLoRA as the baseline for efficient knowledge editing.
-GRACE.
-GRACE (Hartvigsen et al., 2023) introduces a discrete key-value codebook approach, where edits are cached as latent space mappings without altering the original model weights.
-This method enables thousands of sequential edits by storing corrections in a retrieval-based codebook, ensuring minimal interference with unrelated inputs and strong locality .
-However, GRACE’s reliance on non-parametric representations limits its ability to generalize beyond memorized edits, as it struggles to integrate new knowledge into the model’s reasoning process.
-WISE.
-WISE (Wang et al., 2024d) uses a dual-memory architecture comprising a main memory for pretrained knowledge and a side memory for editing.
-A routing mechanism decides which memory to use during inference.
-Additionally, knowledge is partitioned into multiple subspaces through sharding to facilitate conflict-free edits, with subsequent merging of the shards into a unified side memory.
-Appendix C Lifelong Editing Results
-In real-world autonomous driving scenarios, sequential knowledge updates are often required.
-We sequentially test the effects of lifelong editing 1, 250, 500, 750, and 1000 times across four baseline methods: Prompt, AdaLora, GRACE, and WISE.
-The lifelong editing results of Qwen2-VL and LLaVA-Onevision are shown in Figure 7 and Figure 8, respectively.
-Regarding Reliability and Generality metrics, the performance of the four baselines shows a gradual decline, albeit with some fluctuations.
-WISE, as a memory-augmented approach, effectively alleviates the knowledge forgetting phenomenon observed in AdaLora, which relies on parameter updates.
-Although the Prompt method demonstrates also effective lifelong editing performance at 250 and 500 times updates, it leds to an Out-of-Memory (OOM) error after 750 times updates.
-Due to the long tokens of multimodal inputs, the codebook struggles to differentiate between distinct representation information, resulting in suboptimal lifelong editing performance for GRACE.
-Methods that preserve the original parameters, such as WISE and GRACE, achieved 100% performance on the Locality metric.
-Due to interference from additional inputs, the Prompt method performs poorly on the Locality metric, with even worse results observed in multimodal locality.
-AdaLoRA employs low-rank matrix-based parameter updates, which minimally affect the original parameters, resulting in a strong performance on the locality metric.
-Appendix D Benchmark Construction Details
-Appendix E Case of ADS-Edit benchmark
+Acknowledgements.
+We would like to express our sincere gratitude to the anonymous reviewers for their thoughtful and constructive feedback.
+This work was supported by the National Natural Science Foundation of China (No.
+62206246, No.
+NSFCU23B2055, No.
+NSFCU19B2027), the Fundamental Research Funds for the Central Universities (226-2023- 00138), Ningbo Natural Science Foundation (2024J020), Yongjiang Talent Introduction Programme (2021A-156-G), and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
+This work was supported by Ant Group and Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph.References
+- (1)
+- Antol et al.
+(2015) Stanislaw Antol, Aishwarya Agrawal, Jiasen Lu, Margaret Mitchell, Dhruv Batra, C.
+Lawrence Zitnick, and Devi Parikh.
+2015.
+VQA: Visual Question Answering.
+In International Conference on Computer Vision (ICCV).
+- Cai and Cao (2024) Yuchen Cai and Ding Cao.
+2024.
+O-Edit: Orthogonal Subspace Editing for Language Model Sequential Editing.
+CoRR abs/2410.11469 (2024).
+doi:10.48550/ARXIV.2410.11469 arXiv:2410.11469
+- Chen (2024) Huajun Chen. 2024. Large Knowledge Model: Perspectives and Challenges. Data Intelligence 3 (2024).
+- Cheng et al.
+(2023) Siyuan Cheng, Bozhong Tian, Qingbin Liu, Xi Chen, Yongheng Wang, Huajun Chen, and Ningyu Zhang.
+2023.
+Can We Edit Multimodal Large Language Models?.
+In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, EMNLP 2023, Singapore, December 6-10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.).
+Association for Computational Linguistics, 13877–13888.
+doi:10.18653/V1/2023.EMNLP-MAIN.856
+- Dettmers et al.
+(2023) Tim Dettmers, Artidoro Pagnoni, Ari Holtzman, and Luke Zettlemoyer.
+2023.
+QLoRA: Efficient Finetuning of Quantized LLMs.
+In Advances in Neural Information Processing Systems 36: Annual Conference on Neural Information Processing Systems 2023, NeurIPS 2023, New Orleans, LA, USA, December 10 - 16, 2023, Alice Oh, Tristan Naumann, Amir Globerson, Kate Saenko, Moritz Hardt, and Sergey Levine (Eds.).
+http://papers.nips.cc/paper_files/paper/2023/hash/1feb87871436031bdc0f2beaa62a049b-Abstract-Conference.html
+- et al (2024) DeepSeek-AI et al. 2024. DeepSeek-V3 Technical Report. arXiv:2412.19437 [cs.CL] https://arxiv.org/abs/2412.19437
+- Fang et al.
+(2024a) Junfeng Fang, Houcheng Jiang, Kun Wang, Yunshan Ma, Xiang Wang, Xiangnan He, and Tat-Seng Chua.
+2024a.
+AlphaEdit: Null-Space Constrained Knowledge Editing for Language Models.
+CoRR abs/2410.02355 (2024).
+doi:10.48550/ARXIV.2410.02355 arXiv:2410.02355
+- Fang et al.
+(2024b) Junfeng Fang, Houcheng Jiang, Kun Wang, Yunshan Ma, Xiang Wang, Xiangnan He, and Tat-Seng Chua.
+2024b.
+AlphaEdit: Null-Space Constrained Knowledge Editing for Language Models.
+CoRR abs/2410.02355 (2024).
+doi:10.48550/ARXIV.2410.02355 arXiv:2410.02355
+- Gupta et al. (2024) Akshat Gupta, Anurag Rao, and Gopala Anumanchipalli. 2024. Model editing at scale leads to gradual and catastrophic forgetting. arXiv preprint arXiv:2401.07453 (2024).
+- Hartvigsen et al.
+(2023) Tom Hartvigsen, Swami Sankaranarayanan, Hamid Palangi, Yoon Kim, and Marzyeh Ghassemi.
+2023.
+Aging with GRACE: Lifelong Model Editing with Discrete Key-Value Adaptors.
+In Advances in Neural Information Processing Systems 36: Annual Conference on Neural Information Processing Systems 2023, NeurIPS 2023, New Orleans, LA, USA, December 10 - 16, 2023, Alice Oh, Tristan Naumann, Amir Globerson, Kate Saenko, Moritz Hardt, and Sergey Levine (Eds.).
+http://papers.nips.cc/paper_files/paper/2023/hash/95b6e2ff961580e03c0a662a63a71812-Abstract-Conference.html
+- Hu et al.
+(2022) Edward J.
+Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, and Weizhu Chen.
+2022.
+LoRA: Low-Rank Adaptation of Large Language Models.
+In The Tenth International Conference on Learning Representations, ICLR 2022, Virtual Event, April 25-29, 2022.
+OpenReview.net.
+https://openreview.net/forum?id=nZeVKeeFYf9
+- Huang et al.
+(2024b) Han Huang, Haitian Zhong, Tao Yu, Qiang Liu, Shu Wu, Liang Wang, and Tieniu Tan.
+2024b.
+VLKEB: A Large Vision-Language Model Knowledge Editing Benchmark.
+In Advances in Neural Information Processing Systems 38: Annual Conference on Neural Information Processing Systems 2024, NeurIPS 2024, Vancouver, BC, Canada, December 10 - 15, 2024, Amir Globersons, Lester Mackey, Danielle Belgrave, Angela Fan, Ulrich Paquet, Jakub M.
+Tomczak, and Cheng Zhang (Eds.).
+http://papers.nips.cc/paper_files/paper/2024/hash/1198b53fa686831d5f0c0860d7ec4f34-Abstract-Datasets_and_Benchmarks_Track.html
+- Huang et al.
+(2024a) Zhijian Huang, Chengjian Feng, Feng Yan, Baihui Xiao, Zequn Jie, Yujie Zhong, Xiaodan Liang, and Lin Ma.
+2024a.
+DriveMM: All-in-One Large Multimodal Model for Autonomous Driving.
+CoRR abs/2412.07689 (2024).
+doi:10.48550/ARXIV.2412.07689 arXiv:2412.07689
+- Jiang et al.
+(2025) Houcheng Jiang, Junfeng Fang, Ningyu Zhang, Guojun Ma, Mingyang Wan, Xiang Wang, Xiangnan He, and Tat-seng Chua.
+2025.
+AnyEdit: Edit Any Knowledge Encoded in Language Models.
+arXiv preprint arXiv:2502.05628 (2025).
+- Kwiatkowski et al.
+(2019) Tom Kwiatkowski, Jennimaria Palomaki, Olivia Redfield, Michael Collins, Ankur P.
+Parikh, Chris Alberti, Danielle Epstein, Illia Polosukhin, Jacob Devlin, Kenton Lee, Kristina Toutanova, Llion Jones, Matthew Kelcey, Ming-Wei Chang, Andrew M.
+Dai, Jakob Uszkoreit, Quoc Le, and Slav Petrov.
+2019.
+Natural Questions: a Benchmark for Question Answering Research.
+Trans.
+Assoc.
+Comput.
+Linguistics 7 (2019), 452–466.
+doi:10.1162/TACL_A_00276
+- Li et al.
+(2024d) Bo Li, Yuanhan Zhang, Dong Guo, Renrui Zhang, Feng Li, Hao Zhang, Kaichen Zhang, Yanwei Li, Ziwei Liu, and Chunyuan Li.
+2024d.
+LLaVA-OneVision: Easy Visual Task Transfer.
+CoRR abs/2408.03326 (2024).
+doi:10.48550/ARXIV.2408.03326 arXiv:2408.03326
+- Li et al.
+(2024a) Jiaqi Li, Miaozeng Du, Chuanyi Zhang, Yongrui Chen, Nan Hu, Guilin Qi, Haiyun Jiang, Siyuan Cheng, and Bozhong Tian.
+2024a.
+MIKE: A New Benchmark for Fine-grained Multimodal Entity Knowledge Editing.
+In Findings of the Association for Computational Linguistics, ACL 2024, Bangkok, Thailand and virtual meeting, August 11-16, 2024, Lun-Wei Ku, Andre Martins, and Vivek Srikumar (Eds.).
+Association for Computational Linguistics, 5018–5029.
+doi:10.18653/V1/2024.FINDINGS-ACL.298
+- Li et al.
+(2024b) Yanze Li, Wenhua Zhang, Kai Chen, Yanxin Liu, Pengxiang Li, Ruiyuan Gao, Lanqing Hong, Meng Tian, Xinhai Zhao, Zhenguo Li, et al.
+2024b.
+Automated Evaluation of Large Vision-Language Models on Self-driving Corner Cases.
+arXiv preprint arXiv:2404.10595 (2024).
+- Li et al.
+(2024c) Yanze Li, Wenhua Zhang, Kai Chen, Yanxin Liu, Pengxiang Li, Ruiyuan Gao, Lanqing Hong, Meng Tian, Xinhai Zhao, Zhenguo Li, Dit-Yan Yeung, Huchuan Lu, and Xu Jia.
+2024c.
+Automated Evaluation of Large Vision-Language Models on Self-driving Corner Cases.
+CoRR abs/2404.10595 (2024).
+doi:10.48550/ARXIV.2404.10595 arXiv:2404.10595
+- Liu et al. (2024a) Zeyu Leo Liu, Shrey Pandit, Xi Ye, Eunsol Choi, and Greg Durrett. 2024a. Codeupdatearena: Benchmarking knowledge editing on api updates. arXiv preprint arXiv:2407.06249 (2024).
+- Liu et al.
+(2024b) Zeyu Leo Liu, Shrey Pandit, Xi Ye, Eunsol Choi, and Greg Durrett.
+2024b.
+CodeUpdateArena: Benchmarking Knowledge Editing on API Updates.
+CoRR abs/2407.06249 (2024).
+doi:10.48550/ARXIV.2407.06249 arXiv:2407.06249
+- Ma et al.
+(2024) Yaohui Ma, Xiaopeng Hong, Shizhou Zhang, Huiyun Li, Zhilin Zhu, Wei Luo, and Zhiheng Ma.
+2024.
+ComprehendEdit: A Comprehensive Dataset and Evaluation Framework for Multimodal Knowledge Editing.
+CoRR abs/2412.12821 (2024).
+doi:10.48550/ARXIV.2412.12821 arXiv:2412.12821
+- Marcu et al.
+(2024) Ana-Maria Marcu, Long Chen, Jan Hünermann, Alice Karnsund, Benoît Hanotte, Prajwal Chidananda, Saurabh Nair, Vijay Badrinarayanan, Alex Kendall, Jamie Shotton, Elahe Arani, and Oleg Sinavski.
+2024.
+LingoQA: Visual Question Answering for Autonomous Driving.
+In Computer Vision - ECCV 2024 - 18th European Conference, Milan, Italy, September 29-October 4, 2024, Proceedings, Part LXXVII (Lecture Notes in Computer Science, Vol.
+15135), Ales Leonardis, Elisa Ricci, Stefan Roth, Olga Russakovsky, Torsten Sattler, and Gül Varol (Eds.).
+Springer, 252–269.
+doi:10.1007/978-3-031-72980-5_15
+- Meng et al.
+(2022) Kevin Meng, David Bau, Alex Andonian, and Yonatan Belinkov.
+2022.
+Locating and Editing Factual Associations in GPT.
+In Advances in Neural Information Processing Systems 35: Annual Conference on Neural Information Processing Systems 2022, NeurIPS 2022, New Orleans, LA, USA, November 28 - December 9, 2022, Sanmi Koyejo, S.
+Mohamed, A.
+Agarwal, Danielle Belgrave, K.
+Cho, and A.
+Oh (Eds.).
+http://papers.nips.cc/paper_files/paper/2022/hash/6f1d43d5a82a37e89b0665b33bf3a182-Abstract-Conference.html
+- Meng et al.
+(2023) Kevin Meng, Arnab Sen Sharma, Alex J.
+Andonian, Yonatan Belinkov, and David Bau.
+2023.
+Mass-Editing Memory in a Transformer.
+In The Eleventh International Conference on Learning Representations, ICLR 2023, Kigali, Rwanda, May 1-5, 2023.
+OpenReview.net.
+https://openreview.net/forum?id=MkbcAHIYgyS
+- Mitchell et al.
+(2022) Eric Mitchell, Charles Lin, Antoine Bosselut, Chelsea Finn, and Christopher D.
+Manning.
+2022.
+Fast Model Editing at Scale.
+In The Tenth International Conference on Learning Representations, ICLR 2022, Virtual Event, April 25-29, 2022.
+OpenReview.net.
+https://openreview.net/forum?id=0DcZxeWfOPt
+- Qian et al.
+(2024) Tianwen Qian, Jingjing Chen, Linhai Zhuo, Yang Jiao, and Yu-Gang Jiang.
+2024.
+NuScenes-QA: A Multi-Modal Visual Question Answering Benchmark for Autonomous Driving Scenario.
+In Thirty-Eighth AAAI Conference on Artificial Intelligence, AAAI 2024, Thirty-Sixth Conference on Innovative Applications of Artificial Intelligence, IAAI 2024, Fourteenth Symposium on Educational Advances in Artificial Intelligence, EAAI 2014, February 20-27, 2024, Vancouver, Canada, Michael J.
+Wooldridge, Jennifer G.
+Dy, and Sriraam Natarajan (Eds.).
+AAAI Press, 4542–4550.
+doi:10.1609/AAAI.V38I5.28253
+- Sima et al.
+(2024) Chonghao Sima, Katrin Renz, Kashyap Chitta, Li Chen, Hanxue Zhang, Chengen Xie, Jens Beißwenger, Ping Luo, Andreas Geiger, and Hongyang Li.
+2024.
+DriveLM: Driving with Graph Visual Question Answering.
+In Computer Vision - ECCV 2024 - 18th European Conference, Milan, Italy, September 29-October 4, 2024, Proceedings, Part LII (Lecture Notes in Computer Science, Vol.
+15110), Ales Leonardis, Elisa Ricci, Stefan Roth, Olga Russakovsky, Torsten Sattler, and Gül Varol (Eds.).
+Springer, 256–274.
+doi:10.1007/978-3-031-72943-0_15
+- Team (2025) Qwen Team. 2025. Qwen2.5-VL. https://qwenlm.github.io/blog/qwen2.5-vl/
+- Wang et al.
+(2024a) Peng Wang, Shuai Bai, Sinan Tan, Shijie Wang, Zhihao Fan, Jinze Bai, Keqin Chen, Xuejing Liu, Jialin Wang, Wenbin Ge, Yang Fan, Kai Dang, Mengfei Du, Xuancheng Ren, Rui Men, Dayiheng Liu, Chang Zhou, Jingren Zhou, and Junyang Lin.
+2024a.
+Qwen2-VL: Enhancing Vision-Language Model’s Perception of the World at Any Resolution.
+CoRR abs/2409.12191 (2024).
+doi:10.48550/ARXIV.2409.12191 arXiv:2409.12191
+- Wang et al.
+(2024c) Peng Wang, Zexi Li, Ningyu Zhang, Ziwen Xu, Yunzhi Yao, Yong Jiang, Pengjun Xie, Fei Huang, and Huajun Chen.
+2024c.
+WISE: Rethinking the Knowledge Memory for Lifelong Model Editing of Large Language Models.
+In The Thirty-eighth Annual Conference on Neural Information Processing Systems.
+https://openreview.net/forum?id=VJMYOfJVC2
+- Wang et al.
+(2024b) Yiwei Wang, Muhao Chen, Nanyun Peng, and Kai-Wei Chang.
+2024b.
+DeepEdit: Knowledge Editing as Decoding with Constraints.
+CoRR abs/2401.10471 (2024).
+doi:10.48550/ARXIV.2401.10471 arXiv:2401.10471
+- Wu et al.
+(2024) Xiaobao Wu, Liangming Pan, William Yang Wang, and Anh Tuan Luu.
+2024.
+AKEW: Assessing Knowledge Editing in the Wild.
+In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, EMNLP 2024, Miami, FL, USA, November 12-16, 2024, Yaser Al-Onaizan, Mohit Bansal, and Yun-Nung Chen (Eds.).
+Association for Computational Linguistics, 15118–15133.
+https://aclanthology.org/2024.emnlp-main.843
+- Xing et al.
+(2024) Shuo Xing, Chengyuan Qian, Yuping Wang, Hongyuan Hua, Kexin Tian, Yang Zhou, and Zhengzhong Tu.
+2024.
+OpenEMMA: Open-Source Multimodal Model for End-to-End Autonomous Driving.
+CoRR abs/2412.15208 (2024).
+doi:10.48550/ARXIV.2412.15208 arXiv:2412.15208
+- Xu et al.
+(2024) Derong Xu, Ziheng Zhang, Zhihong Zhu, Zhenxi Lin, Qidong Liu, Xian Wu, Tong Xu, Wanyu Wang, Yuyang Ye, Xiangyu Zhao, Enhong Chen, and Yefeng Zheng.
+2024.
+Editing Factual Knowledge and Explanatory Ability of Medical Large Language Models.
+In Proceedings of the 33rd ACM International Conference on Information and Knowledge Management, CIKM 2024, Boise, ID, USA, October 21-25, 2024, Edoardo Serra and Francesca Spezzano (Eds.).
+ACM, 2660–2670.
+doi:10.1145/3627673.3679673
+- Yang et al.
+(2025) Wanli Yang, Fei Sun, Jiajun Tan, Xinyu Ma, Qi Cao, Dawei Yin, Huawei Shen, and Xueqi Cheng.
+2025.
+The Mirage of Model Editing: Revisiting Evaluation in the Wild.
+CoRR abs/2502.11177 (2025).
+doi:10.48550/ARXIV.2502.11177 arXiv:2502.11177
+- Yao et al.
+(2024) Ruoyu Yao, Yubin Wang, Haichao Liu, Rui Yang, Zengqi Peng, Lei Zhu, and Jun Ma.
+2024.
+CALMM-Drive: Confidence-Aware Autonomous Driving with Large Multimodal Model.
+CoRR abs/2412.04209 (2024).
+doi:10.48550/ARXIV.2412.04209 arXiv:2412.04209
+- Yao et al.
+(2025) Yunzhi Yao, Jizhan Fang, Jia-Chen Gu, Ningyu Zhang, Shumin Deng, Huajun Chen, and Nanyun Peng.
+2025.
+CaKE: Circuit-aware Editing Enables Generalizable Knowledge Learners.
+arXiv preprint arXiv:2503.16356 (2025).
+- Yao et al.
+(2023) Yunzhi Yao, Peng Wang, Bozhong Tian, Siyuan Cheng, Zhoubo Li, Shumin Deng, Huajun Chen, and Ningyu Zhang.
+2023.
+Editing Large Language Models: Problems, Methods, and Opportunities.
+In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, EMNLP 2023, Singapore, December 6-10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.).
+Association for Computational Linguistics, 10222–10240.
+doi:10.18653/V1/2023.EMNLP-MAIN.632
+- Youssef et al.
+(2025) Paul Youssef, Zhixue Zhao, Daniel Braun, Jörg Schlötterer, and Christin Seifert.
+2025.
+Position: Editing Large Language Models Poses Serious Safety Risks.
+arXiv preprint arXiv:2502.02958 (2025).
+- Zhang et al.
+(2024c) Junzhe Zhang, Huixuan Zhang, Xunjian Yin, Baizhou Huang, Xu Zhang, Xinyu Hu, and Xiaojun Wan.
+2024c.
+MC-MKE: A Fine-Grained Multimodal Knowledge Editing Benchmark Emphasizing Modality Consistency.
+CoRR abs/2406.13219 (2024).
+doi:10.48550/ARXIV.2406.13219 arXiv:2406.13219
+- Zhang et al.
+(2024a) Ningyu Zhang, Bozhong Tian, Siyuan Cheng, Xiaozhuan Liang, Yi Hu, Kouying Xue, Yanjie Gou, Xi Chen, and Huajun Chen.
+2024a.
+InstructEdit: Instruction-Based Knowledge Editing for Large Language Models.
+In Proceedings of the Thirty-Third International Joint Conference on Artificial Intelligence, IJCAI 2024, Jeju, South Korea, August 3-9, 2024.
+ijcai.org, 6633–6641.
+https://www.ijcai.org/proceedings/2024/733
+- Zhang et al.
+(2024b) Ningyu Zhang, Yunzhi Yao, Bozhong Tian, Peng Wang, Shumin Deng, Mengru Wang, Zekun Xi, Shengyu Mao, Jintian Zhang, Yuansheng Ni, Siyuan Cheng, Ziwen Xu, Xin Xu, Jia-Chen Gu, Yong Jiang, Pengjun Xie, Fei Huang, Lei Liang, Zhiqiang Zhang, Xiaowei Zhu, Jun Zhou, and Huajun Chen.
+2024b.
+A Comprehensive Study of Knowledge Editing for Large Language Models.
+CoRR abs/2401.01286 (2024).
+doi:10.48550/ARXIV.2401.01286 arXiv:2401.01286
+- Zhang et al.
+(2023) Qingru Zhang, Minshuo Chen, Alexander Bukharin, Pengcheng He, Yu Cheng, Weizhu Chen, and Tuo Zhao.
+2023.
+Adaptive Budget Allocation for Parameter-Efficient Fine-Tuning.
+In The Eleventh International Conference on Learning Representations.
+https://openreview.net/forum?id=lq62uWRJjiY
+- Zheng et al.
+(2023) Ce Zheng, Lei Li, Qingxiu Dong, Yuxuan Fan, Zhiyong Wu, Jingjing Xu, and Baobao Chang.
+2023.
+Can We Edit Factual Knowledge by In-Context Learning?.
+In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, EMNLP 2023, Singapore, December 6-10, 2023, Houda Bouamor, Juan Pino, and Kalika Bali (Eds.).
+Association for Computational Linguistics, 4862–4876.
+doi:10.18653/V1/2023.EMNLP-MAIN.296